OpenStudy (anonymous):

a line has a slope of -2 and passes through (8,-9) WHAT IS THE EQAUTION OF THE LINE IN SLOPE INTERCEPT FORM?

OpenStudy (anonymous):

well do the y=mx + c

OpenStudy (anonymous):

y= -9 x = 8 and m= -2...

OpenStudy (anonymous):

i cant find the answer so would it be m=-2

OpenStudy (anonymous):

m= slope...

OpenStudy (anonymous):

y=mx+c we are given slope = m = -2 therefore, y= -2x+c since this line passes through (8,-9) ; x=8,y=-9 must satisfy the equation so , -9 = -2*8 + c => -9 = -16 + c so c = -9+16 = 7 hence the required equation is , y=-2x+7

OpenStudy (anonymous):

thats the answer right